OpenStudy (anonymous):

I NEED HELP!!!! Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = quantity x minus nine divided by quantity x plus five. and g(x) = quantity negative seven x minus eight divided by quantity x minus one.

OpenStudy (anonymous):

@elizabeth710 @rational @raiders88 @Ring @timo86m @TheGaMer007 @Tishockjo @Yondaime @is3535 @omarbirjas @ohohaye @Preetha @Phew @Ashleyisakitty @a1234 @aishanuwang @shamim @SydtheKid913 @Schen421 @Data_LG2 @DoShKa_SyRiA @DARTHVADER2900 @DarronW @flawlessswag @frito&cheeseforever @Gprev @HackberryAbby18 @january123 @JupiterJones @kropot72 @kohai @KaylaIsBae @zepdrix @zviking @zeller012 @xitsaliciaa @Compassionate @clamin @Calvin3 @Vocaroo @BlooooooBlahBLIIIIK @braandooll @bluesassytalk @Nnesha @nessachole02 @minisweet4 @mrsean @mt98 @math_ninja

OpenStudy (anonymous):

Sorry people, I just really need help!!

OpenStudy (anonymous):

Could you possibly take a picture of the equation?

OpenStudy (anonymous):

No it was written out..

myininaya (myininaya):

\[f(x)=\frac{x-9}{x+5} , g(x)=\frac{-7x-8}{x-1}\] is this what you mean?

myininaya (myininaya):

don't think you are going to be able to confirm what you want to with that f and g

Vocaloid (vocaloid):

actually, it's possible, it just requires a lot of algebraic simplification to find f(g(x)), take the expression for g(x) and substitute it in for x in f(x), and after simplifying you should get x do the same thing for g(f(x))

Vocaloid (vocaloid):

please tag me if you need to see the numbers worked out

myininaya (myininaya):

Yes I know it is possible to find f(g(x)) and g(f(x)) but the person will not be able to confirm what they wanted @Vocaloid

myininaya (myininaya):

I think there was a type-o since the question says confirm which means they are indeed inverses but these are not inverses.

Vocaloid (vocaloid):

@myininaya Oh, I see, well in that case, I stand corrected.

myininaya (myininaya):

But the person can reach this conclusion by doing that whole plugin in thing and clearing the compound fractions as you mentioned